A rectangle with sides of length a and a-9 has the same perimeter as a triangle with sides of length a+9, a+7, and a-2. Find the perimeter of the rectangle
Answer by ewatrrr(24785)

 
Hi,

2a + 2(a-9) = a+9 + a+7 + a-2
2a + 2a - 18 = 3a + 14
a = 32
Perimeter of the tr1angle is = 41 + 39 + 30 = 110
CHECKING our answer***rectangle has the same perimeter as a triangle
64 + 46 = 110
Wish You the Best in your Studies.